find the mean of the probability distribution of the random variable X if P(X)=⅒, for X= 1,2,2....,10

1
Expert's answer
2022-03-07T17:15:03-0500

"\mu=1\cdot0.1+2\cdot0.1+3\cdot0.1+4\cdot0.1+5\cdot0.1+"

"+6\cdot0.1+7\cdot0.1+8\cdot0.1+9\cdot0.1+10\cdot0.1="

"=(1+2+3+4+5+6+7+8+9+10)\cdot0.1="

"=55\cdot0.1=5.5"